28​% adults favor the use of unmanned drones by police agencies. Twelve U.S. adults are randomly selected. Find the probability that the number of U.S. adults who favor the use of unmanned drones by police agencies is​ (a) exactly​ three, (b) at least​ four, (c) less than eight.
Answer #1

TOPIC:Binomial distribution.

Deline, p = p (of favoring the use of unnamed) deones by police agencies. = 0.28 (in decimals), n = sample size 12 , Keer, X& we need, (x is at least jour). = P(x+4). = 1- P(x44, = 1 - P (x = 3). = 1- [P(x=9) +P(x=1) + P (x=1) ++ (x=3) 21- (0.72) += 1- [ P(x= 8) + P (X= 9) + 8 (X=10) + P(x = 11) + P (x-12) =1-(12)(0,2098 (0.72) + (3) (0.28) ° (0-72) I + (12) (0.28) 10 (